Brown Cow Dairy uses the aging approach to estimate Bad Debt Expense. The balance of each account receivable is aged on the basis of three time periods as follows: (1) 1–30 days old, $12,800; (2) 31–90 days old, $5,800; and (3) more than 90 days old, $3,800. For each age group, the average loss rate on the amount of the receivable due to uncollectibility is estimated to be (1) 4 percent, (2) 12 percent, and (3) 15 percent, respectively. At December 31 (end of the current year), the Allowance for Doubtful Accounts balance was $880 (credit) before the end-of-period adjusting entry is made.

Required:

  1. Prepare a schedule to estimate an appropriate year-end balance for the Allowance for Doubtful Accounts.

  2. What amount of Bad Debt Expense should be recorded on December 31?

  3. If the unadjusted balance in the Allowance for Doubtful Accounts was a $680 debit balance, what amount of Bad Debt Expense should be recorded on December 31?
Answer #1

a) Schedule

Days Account receivable % Uncollectible $ uncollectible
1-30 days 12800 4% 512
31-90 days 5800 12% 696
Over 90 3800 15% 570
Total 22400 1778

b) Bad debt expense = 1778-880 = 898

c) Bad debt expense = 1778+680 = 2458